n=number
4 less than n/2 = 37 less than 2n.
A. Translate the sentence into an equation.
B. Solve the equation.

Answer by Theo(13342)

n/2 - 4 = 2n - 37
multiply both sides of this equation by 2 to get:
n - 8 = 4n - 74
subtract 4n from both sides of this equation to get:
-3n - 8 = - 74
add 8 to both sides of this equation to get:
-3n = -66
divide both sides of this equation by -3 to get:
n = 22

that's your solution.
when n = 22, your equation of n/2 - 4 = 2n - 37 becomes:
22/2 - 4 = 2*22 - 37 which becomes:
11 - 4 = 44 - 37 which becomes:
7 = 7 which confirms the solution is good.

your solution is n = 22